OpenStudy (anonymous):

slope 7/6,y-intercept: 3 a.y=7/6 x+3 b.y=- 7/6 x+3 c.y=7/6 x-3 d.y=6/7 x+3

OpenStudy (anonymous):

y = mx + c m = 7/6 and c = 3 Put and find the equation..

OpenStudy (ash2326):

A line with slope m and y- intercept c is of the form \[y=mx+c\] Now can you do it?